OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that contains the following points? (-2,-20), (0,-4), (4,-20)

OpenStudy (anonymous):

\[y=ax^2+bx+c\] if \(x=0\) you get \(y=a\times 0^2+b\times 0+c=-4\) so you know \(c=-4\)

OpenStudy (anonymous):

if \(x=-2\) you get \[y=a(-2)^2+b(-2)-4=-20\] or \[4a-2b-4=-20\\ 4a-2b=-16\] this is one equation you can use

OpenStudy (anonymous):

and if \(x=4\) you get (cutting to the chase) \[16a+4b=-16\] solve the system \[4a-2b=-16\\ 16a+4b=-16\] and find \(a\) and \(b\)